OpenStudy (anonymous):

Use a direct proof to show that the sum of two even integers is even. Use the contrapositive (indirect proof).

OpenStudy (perl):

you want a direct proof ?

OpenStudy (perl):

Let a,b be two even integers. then a = 2k , and b = 2m for some integers k,m. a + b = (2k + 2m) = 2 ( k+m) Now a+b is even since a + b = 2 * p , where p = k + m

OpenStudy (anonymous):

contrapositive proof please